OpenStudy (anonymous):

HELP! Simplify completely quantity 7 x squared plus 11 x minus 6 all over quantity 7 x squared minus 10 x plus 3

OpenStudy (jhannybean):

\[\large \cfrac{7x^2+11x-6}{7x^2-10x+3}\] yes?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

quantity 7 x minus 6 over quantity 7 x plus 3 is this correct?

OpenStudy (jhannybean):

Let me check, :P

OpenStudy (anonymous):

okay!

OpenStudy (jhannybean):

No sorry,that isnt :\

OpenStudy (jhannybean):

Let's work through it step by step! :D

OpenStudy (anonymous):

wait! what about this? quantity x plus 2 over quantity 7 x plus 3

OpenStudy (jhannybean):

Starting with the numerator, \(\large 7x^2 +11x -6\), we multiply the leading coefficient wih the constant at the end, \(\large \color{green}7x^2 +11x -\color{green}6\) this helps makeit easier to evaluate our function. We'll have \[\large x^2 +11x -42\]

OpenStudy (jhannybean):

Nope that isnt an option either :\

OpenStudy (jhannybean):

Infact, once wwe write this in simplest form, nothing will be reduced! :D

OpenStudy (anonymous):

dang -_-

OpenStudy (jhannybean):

Can you tell me what two numbers multiply to give -42 and add to give +11?

OpenStudy (jhannybean):

hint: think of your 7's table :P

OpenStudy (anonymous):

I get one or negative when when I do 6 and 7 .

OpenStudy (jhannybean):

well, 7 x ______ = 42?

OpenStudy (anonymous):

\[\frac{ 7x ^{2}+11x-6 }{ 7x ^{2}-10x+3 }=\frac{ 7x ^{2}+14x-3x-6 }{ 7x ^{2}-7x-3x+3 }\] \[=\frac{ 7x \left( x+2 \right)-3\left( x+2 \right) }{ 7x \left( x-1 \right)-3\left( x-1 \right) }\] \[=\frac{ \left( x+2 \right)\left( 7x-3 \right) }{ \left( x-1 \right)\left( 7x-3 \right) }\] \[=\frac{ x+2 }{ x-1 }\

OpenStudy (jhannybean):

we're getting there Surji... haha.

OpenStudy (anonymous):

thank you both! ill take it from here and will explain it t myself!

OpenStudy (jhannybean):

Ok then...

OpenStudy (anonymous):

yw

OpenStudy (jhannybean):

This is my method. \begin{align} \large \cfrac{7x^2+11x-6}{7x^2 -10x+3} &= \large \cfrac{x^2 +11x -42}{x^2 -10x +21} \\ &= \large \cfrac{(x-14)(x+3)}{(x-7)(x-3)} \\ &= \large \cfrac{(x-\cfrac{14}{7})(x+\cfrac{3}{7})}{(x-\cfrac{7}{7})(x-\cfrac{3}{7})} \\ &= \large \cfrac{(x-2)(7x+3)}{(x-1)(7x-3)} \\ \end{align}